OpenStudy (anonymous):

Tammy has a clay vase shaped as a triangular prism with a base area of 42.6 square inches. It is one third filled with water. If there are 304 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth? 14.5 inches 21.4 inches 6.5 inches 7.6 inches

OpenStudy (anonymous):

I just need the formula...

OpenStudy (swissgirl):

http://www.mathsteacher.com.au/year9/ch14_measurement/19_prism/prism.htm is it this shape?

OpenStudy (anonymous):

It didnt show a picture with it..

OpenStudy (swissgirl):

ok well i am assuming that that is the shape so the formula is \( \frac{1} {2} base*height*length = Volume \)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

so the base is 42.6 square inches.... but it doesnt give height or length :(

OpenStudy (anonymous):

oops, its base area not the base soo......

OpenStudy (anonymous):

thanks, I know what formula I needed! (: its V= Bh

OpenStudy (swissgirl):

yupppp

OpenStudy (anonymous):

still not sure what I put for the volume to find the height though...

OpenStudy (swissgirl):

wait let me work it out

OpenStudy (anonymous):

mmk, thank you!!! (:

OpenStudy (swissgirl):

ok got the answer sorry my comp fell and broke :(

OpenStudy (anonymous):

Yay! and dang, that sucks!! :/

OpenStudy (swissgirl):

ok so we know that for a third of the height its 304 cubed So we will just plug this info to our formula but leave out the height since we dont know what a 3rd of the height is 42.6*height=304 height=304/42.6=7.136 Therfore the height for the amount of water in the vase is 7.136 Now that is a third of the height of the vase soooooo our height of the vase is 7.136*3=21.4

OpenStudy (anonymous):

Thank you SOOOO much!! (:

OpenStudy (swissgirl):

np